SUBMILLIMETER-WAVE LINES OF CN^{−}

\underline{TAKAYOSHI AMANO}, Department oof Chemistry, University of Waterloo, Waterloo, ON, Canada N2L 3G1.

High resolution spectroscopy of negative ions is still extremely challenging. Recently Gottlieb et al detected the four low lying rotational lines (J = 1 − 0 to J = 4 − 3) of CN^{−} by using a glow discharge in a mixture of C_{2}N_{2} and Ar or N_{2} a. We extended the measurements to higher frequency up to the J = 7 − 6 transition at 786 GHz. Various types of electric discharges were explored for a better source of negative ions. As shown by Gottlieb et al, a glow discharge yields reasonable signals without extensive signal accumulation. However, it is found that a “hollow anode” discharge with an axial magnetic field of about 250 Gauss is a better source for CN^{−}. The molecular constants have been obtained by including newly measured line frequencies.

aC. A. Gottlieb, S. Brünken, M. C. McCarthy, and P. Thaddeus, J. Chem. Phys. 126, 191101(2007)